# MD 20742, USA

### v2

@@ Abstract.
We study heavy tailed random variables obtained by sampling an observable with a power singularity along an orbit of an expanding map on the circle. In the regime where the observables have infinite variance, we show that the set of limit points of the ergodic sum process, rescaled by $N^{1/s}(\operatorname{ln}N)^{\alpha}$, where $s$ is the index of limiting stable law, equals to the set of piecewise constant increasing functions with finitely many jumps. The proof relies on the multiple Borel-Cantelli Lemma proven in Dolgopyat, Fayad and Liu [J. Mod. Dyn. 18 (2022) 209–289].
@@ 1. Introduction
Let $X_{k}$ be a sequence of independent and identically distributed random variables with mean 0 and variance 1. Let $S_{n}=\sum_{k=1}^{n}X_{k}.$ Consider a function $W_{N}:[0,1]\to\mathbb{R}$ defined by $W_{N}\left(\frac{n}{N}\right)=\frac{S_{n}}{\sqrt{N}}$ at rational points $\frac{n}{N}$ and by linear interpolation in between. The functional central limit theorem states that the distribution of $W_{N}(\cdot)$ converges to the standard Wiener measure as $N\to\infty$. Since Wiener measure has full support, it follows (see e.g. [41]) that with probability $1$ for any continuous function $h$ on $[0,1]$, there is a sequence $N_{j}$ such that $W_{N_{j}}$ converges uniformly to $h.$
It is natural to ask whether one can obtain a smaller set of limits by considering a larger normalization (so that the renormalized function is close to 00 most of the time). Strassen [42] answered this question by showing that if one takes
@@
then with probability $1$, $h$ is a limit point of $\{{\widetilde{W}}_{N}\}$ if and only if $h$ is an absolutely continuous function such that $h(0)=0$ and $\int_{0}^{1}(h^{\prime}(x))^{2}{\rm{d}}x\leq 1.$ This result is referred to as the functional form of law of the iterated logarithm (FLIL) in the literature [16, 39]. FLIL is a generalization of Khintchine-Kolmogoroff’s original form [28, 29] of the law of the iterated logarithm (LIL) and Hartman–Wintner’s LIL [17] for random walks. The method used in [42] for proving FLIL is known as Strassen’s almost sure invariance principle (ASIP). ASIP was further investigated for martingales (see e.g. [16, 20, 43]), which states that a martingale satisfying certain moment conditions is with probability $1$ close to a Brownian motion. Consequently, FLIL holding for Brownian motion is still valid for martingales.
ASIP can be extended to dynamical systems [39] by approximating weakly dependent random variables by martingales. This method has been applied to show ASIP for various dynamical systems (see e.g. [7, 8, 13, 15, 18, 19, 21, 30, 31, 33, 36, 38, 40]). Specifically, these systems include uniformly hyperbolic systems [10], Lorenz attractors [22], non-uniformly hyperbolic systems [32, 34, 35], partially hyperbolic systems [11], dispersing billiards [3, 5] and compact group extensions [14]. However, there are still many examples of dynamical systems where the limit distributions of ergodic sums are non-Gaussian. For these systems, the above approach is not applicable and ASIP does not hold. It is therefore of interest to develop a new method for finding a suitable normalizing sequence $L_{N}$ so that the function ${\tilde{S}}_{N}$ defined by ${\tilde{S}}_{N}(\frac{n}{N})=\frac{S_{n}}{L_{N}}$ has a non-trivial set of limit points.
In this paper, we address the above problem in a simple setting. Let $S_{n}(y)=\sum_{k=1}^{n}\phi(f^{k}y)$ where $f$ is an expanding map of the circle $\mathbb{S}$ and $\phi$ has a power singularity of the type $|y-x|^{-1/s}$ for fixed $x\in\mathbb{S}.$ Assuming that $0<s<2,$ $s\neq 1$ and $x$ is a Lebesgue typical point, we show that a reasonable choice of $L_{N}$ is $L_{N}=N^{1/s}(\operatorname{ln}N)^{\alpha}.$ If $\alpha$ satisfies ${[(r+1)s]}^{-1}<\alpha\leq{(rs)}^{-1},$ then the set of limit points consists of piecewise constant increasing functions starting at 00 with at most $r$ jumps.
@@
The first topic is Laws of Large Numbers for trimmed sums (see [1, 23, 24, 25, 27, 37]). Trimming is an important tool in our analysis since it allows us to discard all the summands in $S_{N}$ except for $r$ largest ones. However, this is only part of our analysis, since the contribution of the retained terms needs to be analyzed in more detail than in the papers mentioned above. We also note that unlike the above mentioned papers, our results apply not only to sums with infinite mean but also to the sums with infinite variance.
The second topic is the application of the classical Borel-Cantelli Lemma to estimate the rate of convergence in the Law of Large Numbers (see [4, 6, 9, 26] and references wherein). This topic is closer to the subject of our paper. However, the papers mentioned above deal with a single $N$ at a time, while we obtain the functional version of the deviation result. Fundamental to our approach is the multiple Borel-Cantelli Lemma proposed recently in [12], which provides a criterion to determine when several rare events occur at the same time scale.
In Section 2, we present the main theorem and sketch its proof using the multiple Borel-Cantelli Lemma. In Section 3, we focus on a key lemma which allows us to control the trimmed sums. The proof in Section 3 uses the reflection principle for the processes under consideration. For completeness, we include some standard proofs as appendices.
@@ 2. An analogue of LIL for Smooth expanding maps
Let $f$ be a $C^{2}$ expanding map of the circle $\mathbb{S}.$ Here we say $f$ is expanding if there exists $\lambda>1$ such that $|f^{\prime}(x)|>\lambda$ for any $x\in\mathbb{S}.$ Then there is a unique $f$-invariant measure $\mu$ that is equivalent to Lebesgue measure [44]. Fix $x,$ and let $\phi$ be an unbounded function of the form
$\phi(y)=a|y-x|^{-1/s}+\psi(y),$
where $\psi$ is a $C^{1}$ function and $s>0.$ Let
$X_{n}(y)=\phi(f^{n}y).$
Denote $\|\cdot\|_{\infty}$ the uniform norm and $L=\|\psi\|_{\infty}.$ Then $\mu(X_{n}>t)=\mu(\phi(y)>t)$ and
(2.1) $\mu\left(|y-x|<\left(\frac{a}{t+L}\right)^{s}\right)\leq\mu(\phi(y)>t)\leq\mu\left(|y-x|\leq\left(\frac{a}{t-L}\right)^{s}\right)$
for $t>L.$ It follows that
(2.2) $\mu(X_{n}>t)\sim a_{s}t^{-s}$
for sufficiently large $t$ and $a_{s}=\rho_{0}a^{s}$, where $\rho_{0}$ is the density of $\mu$ at $x.$ For a positive $\alpha$, set
$W_{N}(t)=\frac{\left(\sum_{n\leq Nt}X_{n}\right)-a_{N}}{N^{1/s}(\operatorname{ln}N)^{\alpha}},$
where $a_{N}=0$ if $s\leq 1$ and $a_{N}=Nt\int\phi(y){\rm{d}}\mu(y)$ if $s>1.$
We regard $W_{N}(\cdot)$ as an element of the space ${\mathbb{D}}[0,1]$ of functions from $[0,1]$ to [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] that are right continuous and have left limits, where we endow [,+][$\infty$,$\infty$][ - ∞ , + ∞ ] with the distance $d(z_{1},z_{2})=\left|\tan^{-1}(z_{1})-\tan^{-1}(z_{2})\right|.$ There are several possible topologies on ${\mathbb{D}}[0,1]$, the most relevant for us is Skorokhod’s $J_{1}$ topology. Let $\Lambda$ be the space of order preserving homeomorphisms on $[0,1].$ Then $J_{1}$ topology is defined using the distance
$d_{J_{1}}(h_{1},h_{2})=\inf_{\lambda\in\Lambda}\max\left\{\|\lambda-Id\|_{\infty},\,\max_{t\in[0,1]}d(h_{1}(t),h_{2}(\lambda(t)))\right\}.$
Two functions are $J_{1}$ close, if their graphs are close with respect to the Hausdorff distance.
In our analysis we shall use several facts about $J_{1}$ topology. First, if two functions are close with respect to the supremum norm, then they are $J_{1}$ close. Second, if one takes ${\mathbb{H}}_{r}$ to be the space of step functions with (at most) $r$ jumps, then $W\in{\mathbb{H}}_{r}$ can be written in the form
@@
where $\bar{r}\leq r$ and $\{c_{m}\}_{m=1}^{r}$ and $\{t_{m}\}_{m=1}^{r}$ are $\bar{r}$–tuples such that $c_{m}$’s are positive and $t_{m}$’s are increasing. Two elements $W_{1}$ and $W_{2}$ of ${\mathbb{H}}_{r}$ with exactly $r$ jumps are $J_{1}$ close provided that $\{c_{m}^{1}\}_{m=1}^{r}$ is close to $\{c_{m}^{2}\}_{m=1}^{r}$ and $\{t_{m}^{1}\}_{m=1}^{r}$ is close to $\{t_{m}^{2}\}_{m=1}^{r}$. Third, $J_{1}$ topology allows us to ignore small jumps.
We also need the following definition. We call $x$ Diophantine if there exists $\epsilon=\epsilon(x)>0$ and $\rho_{0}=\rho_{0}(x)>0$ such that for all positive $\rho\leq\rho_{0}$ and all positive integer $k\leq\epsilon|\operatorname{ln}\rho|,$ $f^{-k}B(x,\rho)\cap B(x,\rho)=\emptyset.$ Due to Lemma 5 in [2], $\mu-$almost every $x$ is Diophantine.
@@
Suppose that either
(i) $10<s<$ 1 and $x$ is non-periodic or
(ii) $21\leq s<$ 2 and $x$ is Diophantine.
Let $\alpha$ be such that $\frac{1}{(r+1)s}<\alpha\leq\frac{1}{rs}$ for some $r\in{\mathbb{N}}.$ Then for almost every $y$ the following holds: the set of $J_{1}$ limit points of $\{W_{N}(t)\}_{N=1}^{\infty}$ is equal to ${\mathbb{H}}_{r}.$
@@
The proof of Theorem 2.1 requires control on the number of times $k\leq n$ such that the events $\left\{y:\,c_{j}-\varepsilon\leq X_{k}(y)\leq c_{j}+\varepsilon\right\}$ occur for some $c_{j}>0$ and $0<\varepsilon<c_{j}.$ If several such events occur for some $n$ we also need to control the occurrence time of these events. To this end, we shall use a generalization of the multiple Borel-Cantelli Lemma proven in [12], those statement is recalled in this subsection.
Consider a sequence $\{\rho_{n}\}_{n=1}^{\infty}$ and a family of events $\{\Omega_{\rho_{n}}\}_{n=1}^{\infty}.$ Suppose that there exist ${\hat{\varepsilon}}>0$ and a decomposition of $\Omega_{\rho_{n}}$ for each $\rho_{n}$ such that
@@
Fix $r\in\mathbb{N^{*}},$ for each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$ we will characterize whether with probability 1, for infinitely many $n,$ the events $\Omega_{\rho_{n},i_{j}}^{k_{j}}$ happen for all $1\leq j\leq r.$
For an arbitrary $r$-tuple $0<k_{1}<\dots<k_{r}\leq n,$ we take the separation indices
$\operatorname{Sep}_{n}(k_{1},\dots,k_{r})=\text{Card}\left\{j\in\{0,\dots r-1\}:k_{j+1}-k_{j}\geq s(n)\right\},\quad k_{0}:=0;$
$\widehat{{\rm Sep}}_{n}(k_{1},\dots,k_{r})=\text{Card}\left\{j\in\{0,\dots r-1\}:k_{j+1}-k_{j}\geq\hat{s}(n)\right\},\quad k_{0}:=0$
for functions $s:{\mathbb{N}}\rightarrow{\mathbb{N}},$ $s(n)\leq(\operatorname{ln}n)^{2}$ and $\hat{s}:{\mathbb{N}}\rightarrow{\mathbb{N}},$ $\varepsilon n\leq\hat{s}(n)<n(1-q)/{(2r)}$ for some $0<q<1,$ $0<\varepsilon<(1-q)/{(2r)},$ respectively.
Suppose the function $\sigma.$ and $\sigma_{\cdot,i}$ defined in (2.3) are increasing and there is a sequence ${\varepsilon}_{n}\to 0$ satisfying some of the mixing conditions below.
@@
    If $0<k_{1}<k_{2}<\dots<k_{r}\leq n$ are such that $\operatorname{Sep}_{n}(k_{1},\dots,k_{r})=r$ and for each $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$
    $\left(\prod_{j=1}^{r}\sigma_{\rho_{n},i_{j}}\right)(1-{\varepsilon}_{n})\leq\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{n},i_{j}}\right)\leq\left(\prod_{j=1}^{r}\sigma_{\rho_{n},i_{j}}\right)(1+{\varepsilon}_{n});$
@@
    There exists $K>0$ such that if $0<k_{1}<k_{2}<\cdots<k_{r}\leq n$ are such that $\operatorname{Sep}_{n}(k_{1},\dots,k_{r})=m<r$, then
    $\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{n}}\right)\leq{\frac{K\sigma_{{\rho}_{n}}^{m}}{(\operatorname{ln}n)^{100r}}};$
@@
    If $0<k_{1}<\dots<k_{r}<l_{1}<\dots<l_{r}$ are such that $2^{i}<k_{\alpha}\leq 2^{i+1},2^{j}<l_{\beta}\leq 2^{j+1}$ for $1\leq\alpha,\beta\leq r$, $j-i\geq b$ for some constant $b\geq 2,$
@@
    $\mu\left(\left[\bigcap_{\alpha=1}^{r}\Omega^{k_{\alpha}}_{{\rho}_{2^{i}},i_{\alpha}}\right]\bigcap\left[\bigcap_{\beta=1}^{r}\Omega^{l_{\beta}}_{{\rho}_{2^{j}},j_{\beta}}\right]\right)\leq\left(\prod_{\alpha=1}^{r}\sigma_{\rho_{2^{i}},i_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\sigma_{\rho_{2^{j}},j_{\beta}}\right)(1+{\varepsilon}_{i}).$
We consider the events
$A_{*}^{n}=\{\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq n,\,\,\operatorname{Sep}(k_{1},\dots k_{r})=r,\,\,\,\Omega^{k_{j}}_{{\rho}_{n}}\text{ happens}\}$
@@
$A_{I_{1},\dots,I_{r},i_{1},\dots i_{r}}^{n}=\{\exists\,k_{j},\,\,k_{j}/n\in I_{j},\,\Omega^{k_{j}}_{{\rho}_{n},i_{j}}\text{ happens}\}.$
We also let
$(M4)_{*}\quad\quad\lim_{n\to\infty}\frac{\mu(A_{*}^{n})}{n^{r}\sigma_{\rho_{n}}^{r}}=1.$
$(M4)_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}\quad\quad\lim_{n\to\infty}\frac{\mu(A_{I_{1},\dots,I_{r},i_{1},\dots,i_{r}}^{n})}{n^{r}\Pi_{j=1}^{r}\left(\sigma_{\rho_{n},i_{j}}|I_{j}|\right)}=1.$
@@
Given a family of events $\{\Omega_{\rho_{n}}\}_{\rho_{n}>0}$ and their decomposition $\Omega_{\rho_{n}}=\bigcup_{i=1}^{p}\Omega_{\rho_{n},i}$, define
@@
We assume that $\rho_{n_{1}}\leq\rho_{n_{2}}$ for $n_{1}\geq n_{2}.$
Suppose ${\mathbf{S}}_{r}<\infty$ and $\Omega_{{\rho}_{n_{1}}}\subset\Omega_{{\rho}_{n_{2}}}$ for $n_{1}\geq n_{2}:$
@@
(a) If $(M1)_{r}$ and $(M2)_{r}$ are satisfied, then with probability $1$, we have that for large $n,$ the events $\Omega_{\rho_{n}}^{k}$ happen for at most $(r-1)$ times.

### v3

@@ Abstract.
We establish functional limit theorems for ergodic sums of observables with power singularities for expanding circle maps. In the regime where the observables have infinite variance, we show that when rescaled by $N^{1/s}(\operatorname{ln}N)^{\alpha}$, the partial sum process has limit points consisting precisely of increasing piecewise constant functions with finitely many jumps. Our approach combines trimming techniques with a multiple Borel-Cantelli argument. It provides a functional law of the iterated logarithm for heavy-tailed processes where classical almost sure invariance principles do not apply.
@@ 1. Introduction
Let $\{X_{k}\}_{k=1}^{\infty}$ be a sequence of independent and identically distributed random variables with mean 0 and variance 1. Define the partial sum $S_{n}=\sum_{k=1}^{n}X_{k}$ and consider the function $W_{N}:[0,1]\to\mathbb{R}$ given by
$W_{N}\left(\frac{n}{N}\right)=\frac{S_{n}}{\sqrt{N}}$
at rational points $n/N,$ extended by linear interpolation elsewhere. The functional central limit theorem states that $W_{N}(\cdot)$ converges in distribution to the standard Wiener measure as $N\to\infty$. Since Wiener measure has full support on $C([0,1])$, it follows that almost surely, for any continuous $h\in C([0,1])$, there exists a subsequence $N_{j}$ such that $W_{N_{j}}$ converges uniformly to $h$ (see e.g. [28]).
It is natural to ask whether one can obtain a smaller set of limits by considering a larger normalization. Strassen [29] answered this question by defining
@@
and proving that almost surely, the limit points of $\{\widetilde{W}_{N}\}$ are precisely the absolutely continuous functions $h$ with $h(0)=0$ and $\int_{0}^{1}|h^{\prime}(x)|^{2}{\rm{d}}x\leq 1.$ This result is referred to as the functional form of law of the iterated logarithm (FLIL) in the literature [12, 27]. FLIL generalizes the classical LIL for random variables [13, 21, 22].
Strassen’s proof relies on the almost sure invariance principle (ASIP), which establishes that, under suitable conditions, a martingale is almost surely close to a Brownian motion. Consequently, FLIL extends to martingales satisfying appropriate moment conditions [12, 30]. The ASIP framework further extends to dynamical systems via martingale approximations for weakly dependent processes [27], establishing FLIL for systems including uniformly hyperbolic systems [8], Lorenz attractors [15], non-uniformly hyperbolic systems [14, 23, 24, 25], partially hyperbolic systems [9], dispersing billiards [3, 5] and compact group extensions [11].
However, there are still many examples of dynamical systems where the limit distributions of ergodic sums are non-Gaussian. For these systems, the aforementioned approach is not applicable, and ASIP does not hold. It is therefore of interest to develop a new method for determining a suitable normalizing sequence $L_{N}$ such that the function
${\tilde{S}}_{N}\left(\frac{n}{N}\right)=\frac{S_{n}}{L_{N}}$
has a non-trivial set of limit points.
In this paper, we address the above problem in a simple setting. Let
$S_{n}(y)=\sum_{k=1}^{n}\phi(f^{k}y),$
where $f$ is an expanding circle map $\mathbb{S},$ and $\phi$ has a power singularity of the form $|y-x|^{-1/s}$ for a fixed point $x\in\mathbb{S}.$ Assuming that $0<s<2$ and $x$ is a Lebesgue typical point, we show that a natural normalizing sequence is
$L_{N}=N^{1/s}(\operatorname{ln}N)^{\alpha}.$
If $\alpha$ satisfies
${[(r+1)s]}^{-1}<\alpha\leq{(rs)}^{-1},$
then the set of limit points consists of piecewise constant increasing functions starting at 0 with at most $r$ jumps.
@@
The first topic is laws of large numbers for trimmed sums (see [1, 17, 18, 20, 26]). Trimming is an important tool in our analysis, as it allows us to discard all summands in $S_{N}$ except for the $r$ largest ones. However, this is only part of our analysis, as the contribution of the retained terms requires more detailed analysis than in the papers mentioned above. Additionally, unlike the studies cited above, our results apply not only to sums with infinite mean but also to sums with infinite variance.
The second topic involves the application of the classical Borel-Cantelli lemma to estimate the rate of convergence in laws of large numbers (see [4, 6, 7, 19] and references wherein). This topic is closer to the subject of our paper. However, while the papers mentioned above deal with a single $N$ at a time, we obtain the functional version of the deviation result. Fundamental to our approach is the multiple Borel-Cantelli lemma proposed recently in [10], which provides a criterion for determining when multiple rare events occur at the same time scale.
The paper is structured as follows. Section 2 states the main theorem and outlines its proof via the multiple Borel-Cantelli lemma. Section 3 proves a key lemma controlling trimmed sums using reflection principles. Standard technical results are deferred to appendices.
@@ 2. An analogue of LIL for Smooth expanding maps
Let $f$ be a $C^{2}$ expanding map of the circle $\mathbb{S}.$ Here $f$ is called expanding if there exists $\gamma>1$ such that $|f^{\prime}(x)|>\gamma$ for all $x\in\mathbb{S}.$ It is known that $f$ admits a unique invariant measure $\mu$ that is absolutely continuous with respect to the Lebesgue measure [31]. Fix $x\in\mathbb{S},$ and consider an unbounded observable $\phi$ of the form
$\phi(y)=\frac{a}{|y-x|^{1/s}}+\psi(y),$
where $\psi$ is a $\psi\in C^{1}(\mathbb{S}),$ $a>0,$ and $s>0.$ Define the sequence $X_{k}(y)=\phi(f^{k}y).$ Then
(2.1) $\mu\left(\left\{y:X_{k}(y)>t\right\}\right)\sim a_{s}t^{-s}$
for sufficiently large $t,$ where $a_{s}=\rho_{0}a^{s}$ and $\rho_{0}$ denotes the density of $\mu$ at $x.$ Now for a positive parameter $\alpha,$ define the process
(2.2) $W_{N}(t)=\frac{\left(\sum_{k\leq Nt}X_{k}\right)-a_{N}}{N^{1/s}(\operatorname{ln}N)^{\alpha}},$
with centering term $a_{N}=0$ if $s\leq 1,$ and $a_{N}=Nt\int\phi(y){\rm{d}}\mu(y)$ if $s>1.$
We regard $W_{N}(\cdot)$ as a sequence of random elements in the space ${\mathbb{D}}[0,1]$ of càdlàg functions from $[0,1]$ to ¯=[$\overline{\mathbb{R}}$$\infty$$\infty$], equipped with the metric
$d(z_{1},z_{2})=\left|\tan^{-1}(z_{1})-\tan^{-1}(z_{2})\right|.$
While several topologies exist on ${\mathbb{D}}[0,1]$, the Skorokhod $J_{1}$ topology is most relevant for our analysis. Let $\Lambda$ denote the set of strictly increasing continuous functions $\lambda:[0,1]\to[0,1]$ with $\lambda(0)=0$ and $\lambda(1)=1$. The $J_{1}$ topology is defined [16] using the distance
$d_{J_{1}}(h_{1},h_{2})=\inf_{\lambda\in\Lambda}\left\{\|h_{1}\circ\lambda-h_{2}\|_{\infty}+\|\lambda-id\|_{\infty}\right\},$
where $\|\cdot\|_{\infty}$ denotes the uniform norm.
Key properties of the $J_{1}$ topology that are employed in our analysis include:
  1. (i)
    Convergence in uniform norm implies convergence in $J_{1}$ topology.
  2. (ii)
    Let ${\mathbb{H}}_{r}$ be the space of step functions with at most $r$ positive jumps, then $W\in{\mathbb{H}}_{r}$ can be written as
@@
    where $\bar{r}\leq r$, the coefficients $\{c_{m}\}_{m=1}^{\bar{r}}$ represent positive amplitudes, and the points $\{t_{m}\}_{m=1}^{\bar{r}}$ are strictly increasing jump times. Two elements $W_{1},W_{2}\in{\mathbb{H}}_{r}$ with exactly $r$ jumps are $J_{1}$-close if their respective amplitudes and jump times are close.
  3. (iii)
    The $J_{1}$ topology allows us to ignore small jumps in the limit.
To present our results, we introduce the notion of Diophantine points. A point $x$ is Diophantine if there exist constants $\epsilon=\epsilon(x)>0$ and $\rho_{0}=\rho_{0}(x)>0$ such that for all $\rho\leq\rho_{0}$ and all positive integers $k\leq\epsilon|\operatorname{ln}\rho|,$
$f^{-k}B(x,\rho)\cap B(x,\rho)=\emptyset.$
This notion of Diophantine points characterizes a specific type of slow recurrence behavior. Rooted in Diophantine approximation, which restricts how well real numbers can be approximated by rational ones, this property effectively prevents dense orbit clustering in dynamical systems. Unlike periodic points, Diophantine points are quantitatively non-periodic. Furthermore, as stated in Lemma 5 of [2], $\mu$-almost every $x$ is Diophantine.
With this definition established, we now proceed to formally state our main theorem.
@@
Under the above hypotheses, assume that either
(i) $0<s<1,$ or
(ii) $1\leq s<2$ and $x$ is Diophantine.
Let $\alpha$ satisfy
$\frac{1}{(r+1)s}<\alpha\leq\frac{1}{rs}$
for some $r\in{\mathbb{N}}.$ Then for almost every $y,$ the following holds: the set of $J_{1}$ limit points of $\{W_{N}(t)\}_{N=1}^{\infty}$ is equal to ${\mathbb{H}}_{r}.$
@@
Given sequences of positive numbers $\{\rho_{N}\}$ and $\{c_{j}\},$ and some $\varepsilon<c_{j}.$ The proof of Theorem 2.1 requires controlling the number of times $k\leq N$ such that the following events occur:
$\left\{y:X_{k}(y)>\rho_{N}\right\}$
and
$\left\{y:\,c_{j}-\varepsilon\leq X_{k}(y)\leq c_{j}+\varepsilon\right\}.$
Furthermore, if multiple such events occur for some $k,$ controlling their timing is essential. To address this,we employ an extension of the multiple Borel–Cantelli lemma, generalizing Theorem 2.12 in [10].
Throughout, we denotes $C$ a constant depending on the fixed number $r$ of the occurring events, which may vary across lines but remains independent of the time scale $N,$ the events $\Omega_{\rho_{N}},$ and the iteration order of $f,$ etc.
Consider a probability space $(\Omega,\mathcal{F},\mu)$ and a family of events $\{\Omega_{\rho_{N}}\}_{N=1}^{\infty}.$ Suppose there exists a constant $\hat{\varepsilon}>0$ and a decomposition of $\Omega_{\rho_{N}}$ such that
@@
For a fixed $r\in\mathbb{N^{*}},$ and each tuple $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$ we characterize whether, with probability 1, the events $\left\{\Omega_{\rho_{N},i_{j}}^{k_{j}}\right\}$ occur for all $1\leq j\leq r$ for infinitely many $N.$
For an $r$-tuple $0<k_{1}<\dots<k_{r}\leq N,$ we take the separation indices
$\operatorname{Sep}_{N}(k_{1},\dots,k_{r})=\sharp\left\{j:0\leq j\leq r-1,\,k_{j+1}-k_{j}\geq s(N)\right\},\,\,\,k_{0}:=0,$
$\widehat{{\rm Sep}}_{N}(k_{1},\dots,k_{r})=\sharp\left\{j:0\leq j\leq r-1,\,k_{j+1}-k_{j}\geq\hat{s}(N)\right\},\,\,\,k_{0}:=0,$
where $s:\mathbb{N}\to\mathbb{N}$ satisfies
$s(N)\leq C(\operatorname{ln}N)^{2},$
and $\hat{s}$$\mathbb{N}$$\rightarrow{\mathbb{N}}$ satisfies
$\varepsilon N\leq\hat{s}(N)<qN/{(2r)}$
for some constants $0<q<1$ and $0<\varepsilon<q/{(2r)}.$
Assume there exists a sequence ${\varepsilon}_{N}\to 0$ satisfying the following conditions, which quantify mixing and non-clustering for $\left\{\Omega^{k}_{{\rho}_{N}}\right\}$ and $\left\{\Omega^{k}_{{\rho}_{N},i}\right\}$ under appropriate separations.
@@
    For $0<k_{1}<k_{2}<\dots<k_{r}\leq N$ with $\operatorname{Sep}_{N}(k_{1},\dots,k_{r})=r$ and all $(i_{1},\dots,i_{r})\in\{1,\dots,p\}^{r},$
    $\left(\prod_{j=1}^{r}\sigma_{\rho_{N},i_{j}}\right)\left(1-{\varepsilon}_{N}\right)\leq\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right)\leq\left(\prod_{j=1}^{r}\sigma_{\rho_{N},i_{j}}\right)\left(1+{\varepsilon}_{N}\right).$
@@
    For $0<k_{1}<k_{2}<\cdots<k_{r}\leq N$ with $\operatorname{Sep}_{N}(k_{1},\dots,k_{r})=m<r$,
    $\mu\left(\bigcap_{j=1}^{r}\Omega^{k_{j}}_{{\rho}_{N}}\right)\leq{\frac{C\sigma_{{\rho}_{N}}^{m}}{(\operatorname{ln}N)^{100r}}}.$
@@
    For $0<k_{1}<\dots<k_{r}<l_{1}<\dots<l_{r}$ where $2^{i}<k_{\alpha}\leq 2^{i+1},$ $2^{j}<l_{\beta}\leq 2^{j+1}$ ($1\leq\alpha,\beta\leq r$), $j-i\geq b$ ($b\geq 2$),
@@
    $\mu\left(\left(\bigcap_{\alpha=1}^{r}\Omega^{k_{\alpha}}_{{\rho}_{2^{i}},i_{\alpha}}\right)\bigcap\left(\bigcap_{\beta=1}^{r}\Omega^{l_{\beta}}_{{\rho}_{2^{j}},j_{\beta}}\right)\right)\leq\left(\prod_{\alpha=1}^{r}\sigma_{\rho_{2^{i}},i_{\alpha}}\right)\left(\prod_{\beta=1}^{r}\sigma_{\rho_{2^{j}},j_{\beta}}\right)\left(1+{\varepsilon}_{N}\right).$
Define the events
$A_{*}^{N}=\left\{\omega\in\Omega:\,\exists\,0<k_{1}<k_{2}<\dots<k_{r}\leq N\,\,\text{such that}\,\,\operatorname{Sep}_{N}(k_{1},\dots,k_{r})=r\,\,\text{and}\,\,\omega\in\Omega^{k_{j}}_{{\rho}_{N}}\right\},$
$A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}=\left\{\omega\in\Omega:\,\exists\,k_{j}\,\,\text{such that}\,\,k_{j}/N\in I_{j}\,\,\text{and }\,\,\omega\in\Omega^{k_{j}}_{{\rho}_{N},i_{j}}\right\},$
@@
$H_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}=\bigcap_{M=1}^{\infty}\bigcup_{N=M}^{\infty}A_{I_{1},\dots,I_{r};i_{1},\dots i_{r}}^{N}.$
Additionally, assume
$(M4)_{*}\quad\quad\limsup_{N\to\infty}\frac{\mu\left(A_{*}^{N}\right)}{N^{r}\sigma_{\rho_{N}}^{r}}\leq C$
and
$(M4)_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}\quad\quad\lim_{N\to\infty}\frac{\mu\left(A_{I_{1},\dots,I_{r};i_{1},\dots,i_{r}}^{N}\right)}{N^{r}\Pi_{j=1}^{r}\left(\sigma_{\rho_{N},i_{j}}|I_{j}|\right)}=1.$
@@
Given a family of events $\{\Omega_{\rho_{N}}\}_{N=1}^{\infty}$ with decompositions $\Omega_{\rho_{N}}=\bigcup_{i=1}^{p}\Omega_{\rho_{N},i}$, define
@@
Assume $\rho_{N_{1}}\leq\rho_{N_{2}}$ and $\Omega_{{\rho}_{N_{1}}}\subset\Omega_{{\rho}_{N_{2}}}$ for $N_{1}\geq N_{2}.$
Case 1: ${\mathbf{S}}_{r}<\infty.$
  • (a)
@@
    If $(M1)_{r}$ and $(M2)_{r}$ hold, then with probability $1,$ the events $\{\Omega_{\rho_{N}}^{k}\}_{k=1}^{N}$ occur at most (r1)(r-1) times for all sufficiently large $N.$
  • (b)